OpenStudy (anonymous):

what is a derivative for e^x/5

OpenStudy (shivaniits):

it would be \[1/5\times e ^{x/5}\]

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

\[\frac{d}{dx}e^u=e^u \frac{du}{dx} \implies \frac{d}{dx}e^{\frac{x}{5}}=e^{\frac{x}{5}}(\frac{1}{5})=\frac{e^{\frac{x}{5}}}{5}\]